f(x)=8+x
g(x)=x-5
If d(x)=f(x)-g(x), find d(x)

Answer by Nate(3500)

d(x) = f(x) - g(x)
d(x) = (8 + x) - (x - 5)
d(x) = 8 + x - x + 5 = 13